Suppose a handbill publisher can buy a new duplicating machine for CAD875 and the duplicator has a 1-year life. The machine is expected to contribute CAD916 to the year’s net revenue.

2.1 What is the expected rate of return?

2.2 If the real interest rate at which funds can be borrowed to purchase the machine is five and a half percent, will the publisher choose to invest in the machine? Explain.
Answer #1

(2.1)

Expected rate of return (ROR) = (Net revenue / Cost) - 1 = (916 / 875) - 1 = 1.0469 - 1 = 0.0469 = 4.69%

(2.2)

Since expected ROR is lower than cost of borrowing, investment should not be made.